# Two-Phase Channel Estimation for RIS-Assisted THz Systems with Beam Split

### v2

@@ III-B3 Sparse Recovery Reformulation
$\mathbf{H}_{k}^{\mathrm{cas}}[m]=\mathbf{\Xi}_{\mathrm{R},m}\mathbf{X}_{k,m}\mathbf{C}_{\mathrm{T},m}^{\mathsf{H}},$ (18)
where $\mathbf{X}_{k,m}\in\mathbb{C}^{Q_{\mathrm{B}}\times Q_{\mathrm{T}}}$ is the total angle-excluded matrix, each row of $\mathbf{X}_{k,m}$ is a superposition of row subsets in $\bar{\mathbf{X}}_{k,m}\triangleq\bar{\boldsymbol{\beta}}_{k,m}\otimes\bar{\mathbf{\Sigma}}_{m}$, i.e., $\mathbf{X}_{k,m}(i,:)=\sum_{j\in\mathcal{P}_{i}}{\bar{\mathbf{X}}_{k,m}}(j,:)$. $\mathcal{P}_{i}$ denotes the set of indices associated with the columns in $\tilde{\mathbf{\Xi}}_{\mathrm{R},m}$ that match the $i$-th column of $\mathbf{\Xi}_{\mathrm{R},m}$. Therefore, there are at most ${L}_{\mathrm{tot},k}=L\times J_{k}$ nonzero entries in $\mathbf{X}_{k,m}$. Employing the Kronecker product property, the vectorized RIS-assisted cascaded channel $\mathbf{h}_{k}^{\mathrm{cas}}[m]\in\mathbb{C}^{N_{\mathrm{T}}N_{\mathrm{R}}\times Q_{\mathrm{B}}Q_{\mathrm{T}}}$ can be given by
@@
Note that the observation matrix $\mathbf{\Psi}_{m}\in\mathbb{C}^{PT\times Q_{\mathrm{B}}Q_{\mathrm{T}}}$ is a sparse matrix. To map this problem to the real number domain for resolution, we then transform the complex form in (21) to the equivalent real form
@@
By systematically assigning symbols to each matrix in (23), the sparse recovery problem can be reformulated as
@@
For simplicity, we use a simplified subscript notation by omitting the index for the matrices and vectors. By defining $T_{\mathrm{tot}}\triangleq PT$ and $Q_{\mathrm{tot}}\triangleq Q_{\mathrm{B}}Q_{\mathrm{T}}$, let $\boldsymbol{y}\in\mathbb{C}^{T_{\mathrm{tot}}}$ denote the measurement vector, $\mathbf{\Psi}\in\mathbb{C}^{T_{\mathrm{tot}}\times Q_{\mathrm{tot}}}$ denote the observation matrix and $\boldsymbol{x}\in\mathbb{C}^{Q_{\mathrm{tot}}}$ denote the sparse representation vector to be estimated with sparsity ${L}_{\mathrm{tot}}$.
III-B4 Sparse Recovery Algorithm Analysis
@@
In designing our channel estimation algorithm, the selection of an appropriate sparse recovery algorithm is crucial due to the unique sparsity characteristics of RIS-assisted THz channels. We considered LS, OMP, and GAMP for this purpose.
a) LS: The LS algorithm, while straightforward, solves the sparse recovery problem by minimizing the residual sum of squares. However, its computational complexity, which is $\mathcal{O}(Q_{\mathrm{tot}}^{3})$, will become significantly high for large matrix dimensions. Consequently, LS is not practical for the initial full CSI estimation phase due to its high computational demands.
b) OMP: OMP is a widely used greedy algorithm that reconstructs the signal by iteratively selecting the dictionary atoms that best match the current residual. The computational complexity of OMP is $\mathcal{O}(L_{\mathrm{tot}}Q_{\mathrm{tot}}^{2})$, which is relatively efficient. However, OMP has a nature of “hard estimation”, which means that it makes definitive decisions at each step without retaining detailed probabilistic information about the signal. This approach can lead to suboptimal performance, particularly in noisy environments or with complex signal structures.
c) GAMP: In contrast to OMP, GAMP uses a probabilistic approach within a message-passing framework to perform global updates, which will maintain a distribution of possible values for each signal element. The computational complexity of GAMP is $\mathcal{O}(Q_{\mathrm{tot}}^{2})$ which is similar to OMP but involves more sophisticated probabilistic calculations. GAMP’s “soft estimation” process retains probabilistic information throughout the iterations, which leads to more accurate estimation. This makes GAMP particularly advantageous in environments with high noise levels or complex signals.
Despite the similar computational complexity of GAMP and OMP, GAMP’s superior accuracy and robustness make it the preferred choice for initial full CSI estimation [^2] Detailed investigation of CBS-based GAMP scheme has been discussed in our previous work [19]. Additionally, due to the probabilistic model employed by the “soft estimation” process, GAMP can ensure more reliable angular information determination and avoid the errors caused by “hard estimation”.
Moreover, for the final stage of channel estimation involving the remaining subcarriers, we employ the LS algorithm. The angular information obtained from the initial estimation phase significantly reduces the dictionary matrix dimensions, which will make LS computationally feasible and effective. This stage can benefit more from the simplicity and accuracy of LS method.
III-C Common Angle Estimation
@@
In this subsection, we focus on the estimation of common angles. Using the selected GAMP algorithm from Section III-B, we can obtain the sparse signal vector $\boldsymbol{x}$, denoted as $\hat{\mathbf{x}}_{k,m}=\left[x_{1},\cdots,x_{Q_{\mathrm{tot}}}\right]^{\mathsf{T}}$. By substituting $\hat{\mathbf{x}}_{k,m}$ into (19), the estimated cascaded channel can be expressed as
@@
This method implies that we could perform the estimation for each subcarrier independently to obtain the channel estimates for all subcarriers. However, this requires implementing the estimation algorithm $M$ times, which would result in an unacceptably high computational complexity. While the beam split effect complicates the channel estimation process, it also offers some insights. It is noticed in (11) that different subcarriers share common physical directions, with the only different term $\eta_{m}$. This indicates that the physical angles remain constant across different subcarriers. Inspired by this, we perform the cascaded channel estimation on only a small subset of subcarriers to acquire the common physical angles. We select $\hat{\mathbf{h}}_{k}^{\mathrm{cas}}[m]$ for $m\in\{1,M/2\}$. Given this estimation, we can further estimate the common directions, i.e., the DoDs at the BS and the coupled directions at the RIS.
III-C1 Common DoDs Estimation at BS
Given the estimated cascaded channel $\hat{\mathbf{h}}_{k}^{\mathrm{cas}}[m]=\mathbf{\Pi}_{m}\hat{\mathbf{x}}_{k,m},\bar{m}\in\left\{1,M/2\right\}$, it is noticed that the $q$-th ($q=1,\dots,Q_{\mathrm{tot}}$) entry in $\hat{\mathbf{x}}_{k,\bar{m}}$ corresponds to the $q$-th column in $\mathbf{\Pi}_{\bar{m}}(:,q)$. This means that given the accurate estimation of $\hat{\mathbf{x}}$, we only need to determine its non-zero positions corresponding to the columns in the matrix $\mathbf{\Pi}_{\bar{m}}$. Since $\mathbf{\Pi}_{\bar{m}}\triangleq\left(\mathbf{C}_{\mathrm{T},\bar{m}}^{*}\otimes\mathbf{\Xi}_{\mathrm{R},\bar{m}}\right)$ is in one-to-one correspondence with the ARVs, this relationship allows us to directly obtain the DoDs at BS. However, inevitable noise and estimation errors can cause power leakage in the estimated $\hat{\mathbf{x}}_{k,\bar{m}}$, which will lead to more than $\Sigma_{k}$ non-zero entries. To address this, we propose a heuristic method to determine the accurate non-zero positions, referred to as common physical support, and calculate the corresponding physical DoDs at the BS, as summarized in Algorithm 1.
Algorithm 1 DoDs Estimation at BS
@@ III-C1 Common DoDs Estimation at BS
In the proposed method, we first establish the normalized energy basis vector $\tilde{\boldsymbol{\epsilon}}_{\bar{m}}$, where each entry is the normalized value of the corresponding element in $\hat{\mathbf{x}}_{k,\bar{m}}$ (Step $2$). As observed in Section III-B3, $\mathbf{\Pi}_{\bar{m}}$ can be regarded as a block matrix of $Q_{\mathrm{T}}$ blocks, each containing $Q_{\mathrm{B}}$ columns. Since each block in $\mathbf{\Pi}_{\bar{m}}$ is aligned with a physical DoD at BS, $\tilde{\boldsymbol{\epsilon}}_{\bar{m}}$ can be partitioned in the same way. By accumulating the $\ell_{2}$-norm for each block, we form the merged-energy vector $\boldsymbol{\epsilon}$ to establish a one-to-one correspondence with the physical DoDs at BS (Step $3-5$). The indices corresponding to the $L$ largest elements in $\boldsymbol{\epsilon}$ will form the common physical support $\mathcal{I}^{\mathrm{BS}}$ (Step $6$). Finally, the DoDs $\left\{\hat{\varphi_{l}}\right\}_{l=1}^{L}$ can be obtained using the knowledge of $\mathcal{I}^{\mathrm{BS}}$ and the sampling rule of the CBS dictionary (Step $7$).
@@
III-C2 Common Coupled Angle Estimation at RIS
@@
where $\mathbf{A}_{N}^{\dagger}\mathbf{A}_{N}=\mathbf{I}_{L}$, and $\mathbf{U}_{k,\bar{m}}\triangleq\left[\mathbf{u}_{k,\bar{m},1},\cdots,\mathbf{u}_{k,\bar{m},L}\right]\in\mathbb{C}^{N_{\mathrm{R}}\times L}$ denotes the rest-CSI matrix, which is only related to the AoAs and AoDs at RIS. By substituting (7) into (26), the $l$-th column $\mathbf{u}_{k,\bar{m},l}$ can be expressed as
@@
where $\Sigma_{\bar{m},l}$ denotes the $\left(l,l\right)$-th element in $\Sigma_{\bar{m}}$, and $\gamma_{k,l,j}=\psi_{l}-\theta_{k,j}\in\left[-2,2\right]$ represents the coupled angle for $l=1,\dots,L$ and $j=1,\dots,J_{k}$.
By defining $\mathbf{A}_{\mathrm{R},\bar{m}}\left(\boldsymbol{\gamma}_{k,l}\right)\triangleq\left[\mathbf{a}\left(\gamma_{k,l,1}\right),\cdots,\mathbf{a}\left(\gamma_{k,l,J_{k}}\right)\right]\in\mathbb{C}^{N_{\mathrm{R}}\times J_{k}}$ as the coupled ARM, (27) can be rewritten as
@@
where $\boldsymbol{\chi}_{k,\bar{m},l}\triangleq\Sigma_{\bar{m},l}\boldsymbol{\beta}_{k,\bar{m}}\in\mathbb{C}^{J_{k}}$ denotes the amplitude vector with the path gains and phase shifts for each coupled angle.
The expression in (28) represents a typical DoA estimation problem, which can be effectively addressed using the MUSIC algorithm due to its high resolution and robustness against noise [29]. However, there are two key challenges in directly applying the MUSIC algorithm. Firstly, the effective angle range for $\boldsymbol{\gamma}_{k,l,j}$, which is the difference between two angles $\psi_{l}$ and $\theta_{k,j}$, is doubled compared to the classical angle range. Moreover, the relative frequency parameter $\eta_{m}$ introduced by beam split effect further expands this range. This results in an increased number of potential estimated angles, making it difficult to distinguish the true physical directions. Secondly, the non-strict orthogonality of $\boldsymbol{\chi}_{k,\bar{m},l}$ and limited measurements at each SC can cause energy leakage into the noise space, which further complicates the identification of physical directions.
To address these challenges, we combine the information from two selected subcarriers and apply spatial smoothing preprocessing to the projected rest-CSI matrix $\mathbf{U}_{k,\bar{m}}$. Inspired by the smooth-MUSIC algorithm, we propose a novel DS-MUSIC algorithm to accurately estimate the coupled angles, as described in Algorithm 2, where $\mathrm{peaks}\left(\boldsymbol{x},k\right)$, $\mathrm{maxk}\left(\boldsymbol{x},k\right)$ and $\mathrm{mink}\left(\boldsymbol{x},k\right)$ return $k$ largest peaks, $k$ largest elements and $k$ smallest elements of vector $\boldsymbol{x}$, respectively.
We first preprocess the projected signal $\mathbf{u}_{k,\bar{m},l}$ by dividing the RIS array into multiple overlapping subarrays, with each subarray containing $N_{\mathrm{sub}}$ reflecting elements. There are $N_{\mathrm{a}}=N_{\mathrm{R}}-N_{\mathrm{sub}}+1$ subarrays in total, which satisfies $N_{\mathrm{a}}>N_{\mathrm{sub}}$ and $N_{\mathrm{a}}$, $N_{\mathrm{sub}}>J_{k}$. The rest-CSI $\mathbf{u}_{k,\bar{m},l,n}$ processed by the $n$-th subarray can be expressed as
@@
where $\mathbf{u}_{k,\bar{m},l,n}\triangleq\mathbf{u}_{k,\bar{m},l}\left(n:n+N_{\mathrm{sub}}-1\right)$ and $\mathbf{A}_{\mathrm{R},\bar{m},n}\triangleq\mathbf{A}_{\mathrm{R},\bar{m}}\left(:,n:n+N_{\mathrm{sub}}-1\right)$. Therefore, the covariance matrix $\mathbf{R}_{k,\bar{m},l,n}^{\mathrm{sub}}\in\mathbb{C}^{N_{\mathrm{sub}}\times N_{\mathrm{sub}}}$ can be written as
@@
[…]
1:  $\bar{\boldsymbol{\gamma}}=-2\left(1+\boldsymbol{q}/Q_{\mathrm{R}}\right),\boldsymbol{q}=1,\dots,Q_{\mathrm{B}}$.
@@
9:           $\tilde{p}_{q}^{\mathrm{MUSIC}}=\frac{1}{\left\|\left(\mathbf{\Omega}_{k,\bar{m},l}^{\mathrm{N}}\right)^{\mathsf{H}}\xi_{\mathrm{R},\bar{m},q}^{\mathrm{sub}}\right\|_{2}^{2}}$.
@@
where $\tilde{\mathbf{U}}_{k,\bar{m},l}\triangleq\left[\mathbf{u}_{k,\bar{m},l,1},\dots,\mathbf{u}_{k,\bar{m},l,N_{\mathrm{a}}}\right]\in\mathbb{C}^{N_{\mathrm{sub}}\times N_{\mathrm{a}}}$ is the concatenation of the subspace-supported vectors $\mathbf{u}_{k,\bar{m},l,n}$ as defined in (29). Then, the eigendecomposition of $\mathbf{R}_{k,\bar{m},l}^{u}$ can be expressed as

### v3

@@ III-C Sparse Recovery Reformulation
$\mathbf{H}_{k}^{\mathrm{cas}}[m]=\mathbf{\Xi}_{\mathrm{R},m}\mathbf{X}_{k,m}\mathbf{C}_{\mathrm{T},m}^{\mathsf{H}}.$ (16)
Recalling (14) that the simplified dictionary $\mathbf{\Xi}_{\mathrm{R},m}$ is constructed by removing duplicate columns from the original dictionary $\tilde{\mathbf{\Xi}}_{\mathrm{R},m}$, the corresponding angle excluded matrix $\mathbf{X}_{k,m}\in\mathbb{C}^{Q_{\mathrm{B}}\times Q_{\mathrm{T}}}$ is the merged version of $\bar{\mathbf{X}}_{k,m}\in\mathbb{C}^{Q_{\mathrm{B}}\times Q_{\mathrm{T}}}$. Specifically, each row of $\mathbf{X}_{k,m}$ is a superposition of row subsets in $\bar{\mathbf{X}}_{k,m}$, i.e., $\mathbf{X}_{k,m}(i,:)=\sum_{j\in\mathcal{P}_{i}}{\bar{\mathbf{X}}_{k,m}}(j,:)$. $\mathcal{P}_{i}$ denotes the set of indices associated with the columns in $\tilde{\mathbf{\Xi}}_{\mathrm{R},m}$ that match the $i$-th column of $\mathbf{\Xi}_{\mathrm{R},m}$. In the cascaded channel, the equivalent angles are the combination of DoDs at BS and coupled angles at RIS, which may overlap and result in fewer effective paths. Therefore, there are ${L}_{\mathrm{tot}}\leq L\times J_{k}$ non-zero entries in $\mathbf{X}_{k,m}$. Employing the Kronecker product property, the vectorized RIS-assisted cascaded channel $\mathbf{h}_{k}^{\mathrm{cas}}[m]\in\mathbb{C}^{N_{\mathrm{T}}N_{\mathrm{R}}\times Q_{\mathrm{B}}Q_{\mathrm{T}}}$ can be given by
@@
Note that the observation matrix $\mathbf{\Psi}_{m}\in\mathbb{C}^{PT\times Q_{\mathrm{B}}Q_{\mathrm{T}}}$ is a sparse matrix.
To facilitate the application of following sparse recovery algorithms and improve numerical stability [32], we transform the complex form in (19) to the equivalent real form
@@
where ${}\Re\{\cdot\}$ and ${}\Im\{\cdot\}$ represent the real and imaginary parts, respectively. By systematically assigning symbols to each matrix in (21), the sparse recovery problem can be reformulated as
@@
For simplicity, we adopt a simplified subscript notation by omitting the indices for the matrices and vectors. Let $T_{\mathrm{tot}}\triangleq 2PT$ and $Q_{\mathrm{tot}}\triangleq 2Q_{\mathrm{B}}Q_{\mathrm{T}}$. Define $\boldsymbol{y}\in\mathbb{C}^{T_{\mathrm{tot}}}$, $\mathbf{\Psi}\in\mathbb{C}^{T_{\mathrm{tot}}\times Q_{\mathrm{tot}}}$ and $\boldsymbol{x}\in\mathbb{C}^{Q_{\mathrm{tot}}}$ as the measurement vector, the observation matrix and the sparse representation vector to be estimated with sparsity ${L}_{\mathrm{tot}}$, respectively. This sparse recovery formulation forms the basis for the subsequent channel estimation process. In what follows, we analyze different sparse recovery algorithms and motivate the selection of GAMP as the preferred approach for the cascaded channel estimation.
III-D Sparse Recovery Algorithm Analysis
@@
Selecting an appropriate sparse recovery algorithm is critical for channel estimation in RIS-assisted THz systems due to the unique sparsity characteristics and error propagation effects. This section evaluates three widely used algorithms: LS, OMP, and GAMP.
a) LS: The LS algorithm solves the linear regression problem by minimizing the residual sum of squares. However, it lacks sparsity constraints, which prevents it from effectively capturing the sparse structure of the channel. As a result, LS is unsuitable for the initial full CSI estimation phase, where a sparse representation is required.
b) OMP: OMP is a greedy algorithm that reconstructs the signal by iteratively selecting dictionary atoms that best match the current residual. Since it follows a “hard estimation” approach, it makes definitive decisions at each iteration without retaining probabilistic information. This approach makes OMP highly sensitive to noise and prone to errors in angle estimation. Without a probabilistic model, it cannot compensate for these errors in later processing, which reduces the robustness of the overall estimation framework.
c) GAMP: Unlike OMP, GAMP adopts a probabilistic message-passing framework that maintains a distribution over signal values throughout the estimation process. This “soft estimation” mechanism allows GAMP to incorporate uncertainty in the estimation process and mitigate the impact of noise. Instead of selecting a single best estimate at each step, GAMP provides a probability distribution for each angle estimate, which improves the reliability of angular information. This characteristic is particularly advantageous in THz channel estimation, where accurate CSI acquisition directly affects overall system performance.
In summary, GAMP provides superior accuracy and robustness, which makes it the preferred choice for initial full CSI estimation [^4] A detailed investigation of the CBS-based GAMP scheme is provided in our previous work [19].. Although GAMP can estimate the full channel across all subcarriers, performing this estimation for every subcarrier individually or jointly would lead to substantial computational complexity. To balance estimation accuracy and computational efficiency, we apply GAMP to only two selected subcarriers. The estimated angular parameters are then used to reconstruct the CSI for the remaining SCs.
IV Common Angle Estimation
@@
Utilizing the selected GAMP algorithm from Section III, we can obtain the sparse signal vector $\boldsymbol{x}$, denoted as $\hat{\mathbf{x}}_{k,m}=\left[x_{1},\cdots,x_{Q_{\mathrm{tot}}}\right]^{\mathsf{T}}$. By substituting $\hat{\mathbf{x}}_{k,m}$ into (17), the estimated cascaded channel can be expressed as
@@
This method implies that we could perform the estimation for each SC independently to obtain the channel estimates for all SCs. However, this requires implementing the estimation algorithm $M$ times, which would result in an unacceptably high computational complexity. While the beam split effect complicates the channel estimation process, it is noticed in (11) that different SCs share common physical directions, with the only different term $\eta_{m}$ [20]. This indicates that the physical angles remain constant across different SCs, which we refer to as the common angles. Inspired by this, we perform the cascaded channel estimation on only a small subset of SCs to acquire the common physical angles. We select $\hat{\mathbf{h}}_{k}^{\mathrm{cas}}[m]$ for $m\in\{1,M/2\}$, which will be explained in Section IV-B. Given this estimation, we can further estimate the common directions, i.e., the DoDs at the BS and the coupled directions at the RIS.
IV-A Common DoDs Estimation at BS
Given the estimated cascaded channel $\hat{\mathbf{h}}_{k}^{\mathrm{cas}}[m]=\mathbf{\Pi}_{m}\hat{\mathbf{x}}_{k,m},\bar{m}\in\left\{1,M/2\right\}$, we propose an energy maximum (EnM) algorithm to estimate the common DoDs at BS [^5] The two SCs at positions 1 and $M/2$ are selected due to the requirement of coupled angle estimation at RIS in Section IV-B.. Notice that while the EnM algorithm can operate on any single SC, leveraging CSI from multiple SCs improves estimation accuracy [22]. Therefore, we accumulate the CSI obtained from cascaded channel estimation in Step 1. It is noticed that the $q$-th ($q=1,\dots,Q_{\mathrm{tot}}$) entry in $\hat{\mathbf{x}}_{k,\bar{m}}$ corresponds to the $q$-th column in $\mathbf{\Pi}_{\bar{m}}(:,q)$. This means that given the accurate estimation of $\hat{\mathbf{x}}$, we only need to determine its non-zero positions corresponding to the columns in the matrix $\mathbf{\Pi}_{\bar{m}}$. Since $\mathbf{\Pi}_{\bar{m}}\triangleq\left(\mathbf{C}_{\mathrm{T},\bar{m}}^{*}\otimes\mathbf{\Xi}_{\mathrm{R},\bar{m}}\right)$ is in one-to-one correspondence with the ARVs, this relationship allows us to directly obtain the DoDs at BS. However, inevitable noise and estimation errors can cause power leakage in the estimated $\hat{\mathbf{x}}_{k,\bar{m}}$, which will lead to more than ${L}_{\mathrm{tot}}$ non-zero entries. To address this, we propose a heuristic method to determine the accurate non-zero positions, referred to as common physical support, and calculate the corresponding physical DoDs at the BS, as summarized in Algorithm 1.
Algorithm 1 EnM Based DoDs Estimation at BS
@@ IV-A Common DoDs Estimation at BS
In the proposed method, we first establish the normalized energy basis vector $\tilde{\boldsymbol{\epsilon}}_{\bar{m}}$, where each entry is the normalized value of the corresponding element in $\hat{\mathbf{x}}_{k,\bar{m}}$ (Step $2$). As observed in Section III-C, $\mathbf{\Pi}_{\bar{m}}$ can be regarded as a block matrix of $Q_{\mathrm{T}}$ blocks, each containing $Q_{\mathrm{B}}$ columns. Since each block in $\mathbf{\Pi}_{\bar{m}}$ is aligned with a physical DoD at BS, $\tilde{\boldsymbol{\epsilon}}_{\bar{m}}$ can be partitioned in the same way. By accumulating the $\ell_{2}$-norm for each block, we form the merged-energy vector $\boldsymbol{\epsilon}$ to establish a one-to-one correspondence with the physical DoDs at BS (Step $3-5$). Then, the indices corresponding to the $L$ largest elements in $\boldsymbol{\epsilon}$ will form the common physical support $\mathcal{I}^{\mathrm{BS}}$ (Step $6$), where the function $\mathrm{maxk}\left(\boldsymbol{\epsilon},L\right)$ returns the $L$ largest elements in vector $\boldsymbol{\epsilon}$ Finally, the DoDs $\left\{\hat{\varphi_{l}}\right\}_{l=1}^{L}$ can be obtained using the knowledge of $\mathcal{I}^{\mathrm{BS}}$ and the sampling rule of the CBS dictionary (Step $7$).
@@
IV-B Common Coupled Angle Estimation at RIS
@@
where $\mathbf{A}_{N}^{\dagger}\mathbf{A}_{N}=\mathbf{I}_{L}$, and $\mathbf{U}_{k,\bar{m}}\triangleq\left[\mathbf{u}_{k,\bar{m},1},\cdots,\mathbf{u}_{k,\bar{m},L}\right]\in\mathbb{C}^{N_{\mathrm{R}}\times L}$ denotes the rest-CSI matrix, which is only related to the DoAs and DoDs at RIS. By substituting (7) into (24), the $l$-th column $\mathbf{u}_{k,\bar{m},l}$ can be expressed as
@@
where $\Sigma_{\bar{m},l}$ denotes the $\left(l,l\right)$-th element in $\mathbf{\Sigma}_{\bar{m}}$, and $\gamma_{k,l,j}=\psi_{l}-\vartheta_{k,j}\in\left[-2,2\right]$ represents the coupled angle for $l=1,\dots,L$ and $j=1,\dots,J_{k}$.
Define $\mathbf{A}_{\mathrm{R},\bar{m}}\left(\boldsymbol{\gamma}_{k,l}\right)\triangleq\left[\mathbf{a}_{\mathrm{R},\bar{m}}\left(\gamma_{k,l,1}\right),\cdots,\mathbf{a_{\mathrm{R},\bar{m}}}\left(\gamma_{k,l,J_{k}}\right)\right]\in\mathbb{C}^{N_{\mathrm{R}}\times J_{k}}$ as the coupled ARM, (25) can be rewritten as
@@
where $\boldsymbol{\chi}_{k,\bar{m},l}\triangleq\Sigma_{\bar{m},l}\boldsymbol{\beta}_{k,\bar{m}}\in\mathbb{C}^{J_{k}}$ denotes the amplitude vector with path gains and phase shifts for each coupled angle.
The expression in (26) represents a typical DoA estimation problem, which can be effectively addressed using the MUSIC algorithm due to its high resolution and robustness against noise [33]. However, there are two key challenges in directly applying the MUSIC algorithm. Firstly, the effective angle range for $\boldsymbol{\gamma}_{k,l,j}$, which is the difference between two angles $\psi_{l}$ and $\theta_{k,j}$, is doubled compared to the classical angle range. Moreover, the relative frequency parameter $\eta_{m}$ introduced by beam split effect further expands this range. This results in an increased number of potential estimated angles, making it difficult to distinguish the true physical directions. Secondly, the approximate orthogonality of $\boldsymbol{\chi}_{k,\bar{m},l}$ and limited measurements per SC can cause energy leakage into the noise space, further complicating the identification of physical directions.
To address these challenges, we propose a novel double-sensing MUSIC (DS-MUSIC) algorithm. Firstly, we combine the information from two selected SCs and apply spatial smoothing preprocessing to the projected rest-CSI matrix $\mathbf{U}_{k,\bar{m}}$. Inspired by the spatial smoothing MUSIC algorithm [34], we then introduce the proposed DS-MUSIC algorithm to accurately estimate the coupled angles, as described in Algorithm 2, where $\mathrm{peaks}\left(\boldsymbol{x},k\right)$ and $\mathrm{mink}\left(\boldsymbol{x},k\right)$ return $k$ largest peaks and $k$ smallest elements of vector $\boldsymbol{x}$, respectively.
We first preprocess the projected signal $\mathbf{u}_{k,\bar{m},l}$ by dividing the RIS array into multiple overlapping subarrays, with each subarray containing $N_{\mathrm{sub}}$ reflecting elements. There are $N_{\mathrm{a}}=N_{\mathrm{R}}-N_{\mathrm{sub}}+1$ overlapping subarrays in total. To ensure reliable angle estimation, we enforce the condition $N_{\mathrm{a}}>N_{\mathrm{sub}}$ and $N_{\mathrm{a}},N_{\mathrm{sub}}>J_{k}$ [^6] Increasing $N_{\mathrm{a}}$ provides more snapshots for constructing the sample covariance matrix, which improves its estimation accuracy and the stability of the MUSIC algorithm; on the other hand, a larger $N_{\mathrm{sub}}$ improves the angular resolution by increasing the aperture of each subarray. Since $N_{\mathrm{a}}+N_{\mathrm{sub}}-1=N_{\mathrm{R}}$ is fixed, these parameters must be jointly chosen to balance spatial resolution and covariance estimation accuracy.. The rest-CSI $\mathbf{u}_{k,\bar{m},l,n}$ processed by the $n$-th subarray can be expressed as
@@
where $\mathbf{u}_{k,\bar{m},l,n}\triangleq\mathbf{u}_{k,\bar{m},l}\left(n:n+N_{\mathrm{sub}}-1\right)$ and $\mathbf{A}_{\mathrm{R},\bar{m},n}\triangleq\mathbf{A}_{\mathrm{R},\bar{m}}\left(:,n:n+N_{\mathrm{sub}}-1\right)$. Additionally, $\boldsymbol{\chi}_{k,\bar{m},l}\triangleq\Sigma_{\bar{m},l}\boldsymbol{\beta}_{k,\bar{m}}\in\mathbb{C}^{J_{k}}$ denotes the amplitude vector with the path gains and phase shifts for each coupled angle. Therefore, the covariance matrix $\mathbf{R}_{k,\bar{m},l,n}^{\mathrm{sub}}\in\mathbb{C}^{N_{\mathrm{sub}}\times N_{\mathrm{sub}}}$ can be written as
@@
0:  Projected matrix $\mathbf{U}_{k,\bar{m}}$, the path number $L$ and $J_{k}$ of the BS-RIS and RIS-$k$-th UE channel $\forall k\in\mathcal{K}$, the number of subarrays $N_{\mathrm{a}}$ and quantization level $Q_{\mathrm{R}}$.
1:  $[\bar{\boldsymbol{\gamma}}]_{q}=-2\left(1+q/Q_{\mathrm{R}}\right),q=1,\dots,Q_{\mathrm{B}}$.
@@
9:           $\tilde{\mathbf{p}}_{k,\bar{m},l,q}=\frac{1}{\left\|\left(\mathbf{\Omega}_{k,\bar{m},l}^{\mathrm{N}}\right)^{\mathsf{H}}\xi_{\mathrm{R},\bar{m},q}^{\mathrm{sub}}\right\|_{2}^{2}}$.
@@
where $\tilde{\mathbf{U}}_{k,\bar{m},l}\triangleq\left[\mathbf{u}_{k,\bar{m},l,1},\dots,\mathbf{u}_{k,\bar{m},l,N_{\mathrm{a}}}\right]\in\mathbb{C}^{N_{\mathrm{sub}}\times N_{\mathrm{a}}}$ denotes the concatenation of the subspace-supporting vectors $\mathbf{u}_{k,\bar{m},l,n}$ as defined in (27). The eigendecomposition of $\mathbf{R}_{k,\bar{m},l}^{u}$ can be expressed as